Construction utilizes the Pythagorean Theorem in many instances to insure a “square” structure. In laying out the footers or corners of any building the concept of a 3-4-5 right triangle is applied.

(1)Name two more dimensions of a right triangle that will also yield a square corner and justify.

Answer by oberobic(2304)

You could try: 9, 12, 15.
Check to see if they fit the Pythagorean Theorem's requirements.
C^2 = A^2 + B^2
A = 9
B = 12
C = 15
.
A^2 = 81
B^2 = 144
A^2+B^2 = 81+144
81+144 = 225
sqrt(225) = 15
.
That's correct.
.
Done.